OpenStudy (anonymous):

The number if tickets sold each day for an upcoming performance of Handel's Messiah is given by N(x)=-0.6x^2+13.2x+12, where x is the number of days since the concert was first announced. When will daily ticket sales peak and tickets will be sold that day? Ticket sales will peak ? days after the concert was first announced. The number of tickets sold on that day will be ?

OpenStudy (anonymous):

looks familiar. remember that to find the "vertex" of your parabola you use \[-\frac{b}{2a}\] for the input

OpenStudy (anonymous):

ur trying to find the maximum of the equation

OpenStudy (anonymous):

in your case \[a=-.6,b=13.2\]

OpenStudy (anonymous):

Satellite I am having trouble still with the other I was confused which answer was which can you help me clarify please.

OpenStudy (anonymous):

so your first job is to compute \[-\frac{b}{2a}=-\frac{13.2}{2\times -.6}\]

OpenStudy (anonymous):

no problem. is that much clear? you must compute \[-\frac{b}{2a}\] first to get your answer

OpenStudy (anonymous):

the "leading coefficient" of \[-.6x^2+13.2+12\] is -.6 which is negative so you know you will have a parabola which opens down

OpenStudy (anonymous):

here is a nice picture of it http://www.wolframalpha.com/input/?i=-.6x^2%2B13.2%2B12

OpenStudy (anonymous):

the vertex is the highest point on the parabola. the first coordinate of the vertex will be \[\frac{13.2}{1.2} \] by the formula \[-\frac{b}{2a}\]

OpenStudy (anonymous):

\[\frac{13.2}{1.2}=11\] so the first coordinate of your vertex is 11

OpenStudy (anonymous):

so what do i put in the answer then? 11

OpenStudy (anonymous):

yes that is the first coordinate, correct

OpenStudy (anonymous):

11days

OpenStudy (anonymous):

eleven days. and to find out how many tickets are sold on that day, replace x by 11 and see what you get

OpenStudy (anonymous):

in other words compute \[-.6\times 11^2+13.2\times 11+12\]

OpenStudy (anonymous):

i get 84.6 so i guess you have to round. now i have to run, but don't forget that formula for the vertex. first coordinate will be \[-\frac{b}{2a}\]

OpenStudy (anonymous):

so solve for n?

OpenStudy (anonymous):

just replace x by 11 and you will get \[N(11)=84.6\]

OpenStudy (anonymous):

try it. good luck

OpenStudy (anonymous):

right thx